# MoFO: Momentum-Filtered Optimizer for Mitigating Forgetting in LLM Fine-Tuning

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks. Typically, LLMs are first pre-trained on large corpora and subsequently fine-tuned on task-specific datasets. However, during fine-tuning, LLMs may forget some knowledge acquired in the pre-training stage, leading to a decline in general capabilities. Existing approaches to mitigate forgetting often rely on access to pre-training data, which may be unavailable in many real-world scenarios–such as fine-tuning checkpoint-only open-source LLMs. To address this challenge, we propose a new fine-tuning algorithm termed Momentum-Filtered Optimizer (MoFO). MoFO is an extension of greedy block coordinate descent (BCD) methods: in each iteration, MoFO only updates the model parameters with the largest momentum magnitudes, while keeping all other parameters fixed. MoFO achieves similar fine-tuning performance to the default fine-tuning algorithm while effectively mitigating knowledge forgetting. We validate MoFO through rigorous convergence analysis and extensive experiments, demonstrating its effectiveness in mitigating forgetting without pre-training data.
Machine Learning, ICML
@@ 1 Introduction
The success of large language models (LLMs) lies in their strong capabilities in language understanding and generation. Typically, LLMs are first pre-trained on extensive corpora to acquire general capabilities, then fine-tuned on smaller, task-specific datasets to adapt to particular tasks or domains (Dai & Le, 2015; Kenton & Toutanova, 2019; Radford et al., 2018). However, it has been observed that during the fine-tuning process, LLMs may forget the knowledge acquired in pre-training, leading to a decline in general capabilities (Lin et al., 2023; Chen et al., 2020; Dong et al., 2021; Korbak et al., 2022; Luo et al., 2023). Therefore, addressing the forgetting issue in LLM fine-tuning has become an important research direction.
@@
Given the limitations of these prevalent forgetting-mitigation approaches, we aim to develop an optimization algorithm that neither relies on past data nor introduces additional memory overhead. We adopt an important insight in continual learning (Kirkpatrick et al., 2017): the closer to the previous model, the less forgetting occurs. What optimization method might move a small distance from the initial point, but can still find a minimum [^1] In this paper, we use the term ”minimum” (or ”minima” in the plural) to refer to a parameter configuration whose fine-tuning loss is near its lowest value in a small neighborhood, while acknowledging that this terminology may not strictly represent a local minimum of the fine-tuning loss function. of the loss function? We notice that the classical block coordinate descent (BCD) method (Tseng, 2001) is a good candidate, since it updates only a subset of parameters at each iteration, thus implicitly biased towards closer solutions. Nevertheless, incorporating BCD into LLM fine-tuning presents some challenges. One challenge is that Adam, the predominant optimizer for LLM training (Radford et al., 2018; Zhang et al., 2024c), differs substantially from the optimizer studied in traditional BCD methods (e.g. GD or SGD) (Nutini et al., 2015; Zhao et al., 2014). It complicates both optimizer design and convergence analysis. Consequently, combining BCD with Adam is not a straightforward task.
@@
Due to the page limitation, we provide the review of related works in Appendix A.
@@ 2.1 Motivation
In this work, we conduct exploratory experiments to investigate this correlation among minima produced by different LLM fine-tuning methods. We conduct two sets of experiments. First, we fine-tune Pythia-160M on a subset of the FLAN dataset [^2] The subset used is ‘definite_pronoun_resolution_10templates,’ available at https://huggingface.co/datasets/Muennighoff/flan. The learning rate is 2e-5 and the batch size is set as 64. using Adam (Kingma & Ba, 2014) and Lion (Chen et al., 2024). As illustrated in Figure LABEL:pythia_landscape_lion(a), Adam and Lion converge to distinct minima. Notably, while both optimizers achieve similar fine-tuning losses, the minimum reached by Adam is much closer to the pre-trained model compared to Lion, being only about 20% of Lion’s distance. (Specifically we measure the model distance by the average $L_{2}$ norm difference over parameter blocks; see Appendix D.4 for the formal definition). As shown in Figure LABEL:pythia_landscape_lion(b), Adam exhibits a smaller increase in pre-training loss than Lion. Further, Adam leads to better preservation in general capability than Lion (see Appendix C.1).
Second, we conduct a larger-scale exploratory experiment to test the generality of this observation. We fine-tune a larger model, LLaMA2-7B, on the MetaMathQA dataset (Yu et al., 2024a) using three optimizers: Adam, Lion, and our proposed MoFO optimizer (to be introduced in Section 2.2). To achieve varying distances from the pre-trained model, we fine-tune each model for 0.5, 1, 1.5, and 2 epochs (309 steps per epoch). Details are provided in Appendix D.3. Figure LABEL:relation_fig(a) demonstrates a strong positive correlation between distance and the increase in pre-training loss. Figure LABEL:relation_fig(b) indicates a strong negative correlation with MMLU accuracy (Hendrycks et al., 2021), which measures the preservation of factual knowledge. These findings suggest that in LLM fine-tuning, staying near the pre-trained state can better mitigate forgetting.
@@ 2.2 Formulation of MoFO
Figure 2: Illustration of MoFO.
We formally introduce the Momentum-Filtered Optimizer (MoFO) in Algorithm 1. First, all model parameters are partitioned into $B$ blocks. At each iteration, MoFO first computes the gradient and momentum terms for parameters in each block following the standard rule of Adam, as shown in Lines 5-9. Then, MoFO selects and updates the parameter entries with the largest $\alpha\%$ momentum magnitudes in each parameter block, as shown in Lines 10-13, where the update fraction $\alpha\%$ is a pre-determined hyperparameter. This momentum filtering mechanism is illustrated in Figure 2.
Mathematically, the filter can be represented as follows. Consider a momentum vector $m=(m^{(1)},\dots,m^{(B)})$, where each $m^{(k)}\in\mathbb{R}^{d_{k}}$ corresponds to the $k$-th block of parameters with dimensionality $d_{k}$. The top-$\alpha\%$ filter, denoted as $\texttt{FLT}_{\alpha}(m)$, is defined as $\texttt{FLT}_{\alpha}(m)=(\texttt{FLT}_{\alpha}^{(1)}(m),\dots,\texttt{FLT}_{\alpha}^{(B)}(m))$, where the $i$-th entry of $\texttt{FLT}_{\alpha}^{(k)}(m)$ is given by
$\left[\texttt{FLT}_{\alpha}^{(k)}(m)\right]_{i}$
=$=\begin{cases}1&\text{if }|m^{(k)}_{i}|\text{ is within the top-}\alpha\%\text{ of }|m^{(k)}|\text{ values},\\ 0&\text{otherwise},\end{cases}$
For the parameter partitioning, we note that the network architecture is naturally composed of different modules (e.g., weight matrices, and bias terms). In the PyTorch implementation, the parameters of different modules (along with their gradients and momenta) are naturally stored in separate data tensors. Therefore, we adopt the default partitioning of model parameters as implemented in PyTorch. This allows us to select and update the top $\alpha\%$ parameters in each block without introducing much implementation overhead. See Appendix D.4 for further explanation of the partitioning.
@@
1:  Input: Filtering threshold $\alpha\%$, number of partitions $B$ with the $k$-th partition of size $d_{k}$, hyperparameters $\beta_{1},\beta_{2}$ of Adam optimizer, learning rate schedule $\{\eta_{t}\}$.
@@
11:           $[\texttt{FLT}^{(k)}_{\alpha}(m_{t})]_{i}=1$ if $|(m^{(k)}_{t})_{i}|$ is within the top-$\alpha\%$ of $|m^{(k)}_{t}|$’s values else 0
@@ 3.1 Convergence Result
Suppose that the first and second order momentum hyperparameters $\beta_{1}$ and $\beta_{2}$ satisfy $10<\beta_{1}<\sqrt{\beta_{2}}<$ 1. The learning rate schedule at step $t$ is $\eta_{t}=\eta/\sqrt{t}$ for some $\eta>0$. Then under Assumption 1, MoFO satisfies
@@
$\min_{0\leq t\leq T-1}\|\nabla\mathcal{L}(\theta_{t})\|_{\infty}=\mathcal{O}\left(\frac{\log T}{\sqrt{T}}\right)\quad\text{as }T\to\infty.$
Although MoFO is designed to mitigate forgetting by updating only a small subset of parameters at each step, it is guaranteed to converge to a critical point of the fine-tuning loss function under theoretical assumptions of bounded gradient and Lipschitz smoothness. This result provides theoretical evidence that MoFO can achieve competitive performance in fine-tuning tasks.
Our proof is inspired by the convergence analysis of full-batch Adam from Shi et al. (2021). Suppose that $g_{t}=\nabla\mathcal{L}(\theta_{t-1})$. A pivotal step in their proof involves applying the descent lemma to Adam, resulting in:
$\mathcal{L}(\theta_{t})-\mathcal{L}(\theta_{t-1})\leq-\frac{\eta}{\sqrt{t}}\sum_{i=1}^{d}g_{i,t}\frac{\hat{m}_{i,t}}{\sqrt{\hat{v}_{i,t}}}+\frac{L}{2}\|\theta_{t}-\theta_{t-1}\|_{2}^{2}.$ (1)
Through a series of manipulations, they derive:
$C_{1}\|g_{t}\|_{1}\big{/}\sqrt{t}\leq\mathcal{L}(\theta_{t-1})-\mathcal{L}(\theta_{t})+C_{2}\big{/}t.$ (2)
Summing this inequality from $t=1$ to $T$ yields the convergence result for Adam in terms of a diminishing $\ell_{1}$-norm of gradient, given by
@@
$\min_{1\leq t\leq T}\|g_{t}\|_{1}=\mathcal{O}\left(\frac{\log T}{\sqrt{T}}\right).$ (3)
Stemming from BCD methods, MoFO updates only a subset of parameters at each iteration. Consequently, the summation over all coordinates $i$ in (1) reduces to a summation over the coordinates selected by the momentum filter. Thus, instead of (2), we derive the following inequality:
@@
$C_{1}\|g_{t}\odot\texttt{FLT}_{\alpha}(m_{t})\|_{1}\big{/}\sqrt{t}\leq\mathcal{L}(\theta_{t-1})-\mathcal{L}(\theta_{t})+C_{2}\big{/}t.$ (2’)
We note that the update in MoFO is filtered based on the momentum magnitude rather than the gradient magnitude. This introduces a non-trivial challenge because the left-hand side of (2’) being small does not necessarily imply that the gradient norm is small. This complicates the analysis, and we cannot directly establish a convergence similar to (3).
By analyzing the interaction between momentum magnitudes and gradient norms, we show that the momentum-based filtering in MoFO introduces only a diminishing gap of $\mathcal{O}(1/{\sqrt{t}})$ between $\|g_{t}\odot\texttt{FLT}_{\alpha}(m_{t})\|_{1}$ and the gradient’s $\ell_{\infty}$-norm, $\|g_{t}\|_{\infty}$. Formally, $|g_{t}\|_{\infty}\leq\|g_{t}\odot\texttt{FLT}_{\alpha}(m_{t})\|_{1}+\mathcal{O}(1/{\sqrt{t}})$. The gap is small enough to ensure that the overall convergence rate does not degrade compared to Adam. As a result, we establish a convergence result for MoFO in terms of $\|g_{t}\|_{\infty}$, presented in Theorem 1.

### v4

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks. Typically, LLMs are first pre-trained on large corpora and subsequently fine-tuned on task-specific datasets. However, during fine-tuning, LLMs may forget some knowledge acquired in the pre-training stage, leading to a decline in general capabilities. Existing approaches to mitigate forgetting often rely on access to pre-training data, which may be unavailable in many real-world scenarios—such as fine-tuning checkpoint-only open-source LLMs. To address this challenge, we propose a new fine-tuning algorithm termed Momentum-Filtered Optimizer (MoFO). MoFO is an extension of greedy block coordinate descent (BCD) methods: in each iteration, MoFO only updates the model parameters with the largest momentum magnitudes, while keeping all other parameters fixed. MoFO achieves similar fine-tuning performance to the default fine-tuning algorithm while effectively mitigating knowledge forgetting. We validate MoFO through rigorous convergence analysis and extensive experiments, demonstrating its effectiveness in mitigating forgetting without pre-training data. [^1] Our code is available at https://github.com/YChen-zzz/MoFO.
@@ 1 Introduction
The success of large language models (LLMs) lies in their strong capabilities in language understanding and generation. Typically, LLMs are first pre-trained on extensive corpora to acquire general capabilities, then fine-tuned on smaller, task-specific datasets to adapt to particular tasks or domains (Dai & Le, 2015; Kenton & Toutanova, 2019; Radford et al., 2018). However, it has been observed that during the fine-tuning process, LLMs may forget the knowledge acquired in pre-training, leading to a decline in general capabilities (Lin et al., 2023; Chen et al., 2020; Dong et al., 2021; Korbak et al., 2022; Luo et al., 2023a). Therefore, addressing the forgetting issue in LLM fine-tuning has become an important research direction.
@@
Given the limitations of prevalent forgetting-mitigation approaches, we aim to develop an optimization algorithm that neither relies on past data nor introduces additional memory overhead. In general, our algorithm design goals are twofold: (G1) preservation of pretrained knowledge and (G2) strong performance on the fine-tuning task. To pursue (G1), we adopt a key insight from continual learning (Kirkpatrick et al., 2017): the closer a fine-tuned model stays to the pretrained parameters, the less forgetting tends to occur. At the same time, enforcing proximity too aggressively can hinder adaptation to the new task, thereby harming (G2). This raises a natural question: What optimization method might move a small distance from the initial point, but can still find a minimum [^2] In this paper, we use the term ”minimum” (or ”minima” in the plural) to refer to a parameter configuration whose fine-tuning loss is near its lowest value in a small neighborhood, while acknowledging that this terminology may not strictly represent a local minimum of the fine-tuning loss function. of the loss function?
We notice that the classical block coordinate descent (BCD) method (Tseng, 2001) is a good candidate, since it updates only a subset of parameters at each iteration, thus implicitly biased towards closer solutions. Nevertheless, incorporating BCD into LLM fine-tuning presents some challenges. One challenge is that Adam, the predominant optimizer for LLM training (Radford et al., 2018; Zhang et al., 2024c), differs substantially from the optimizer studied in traditional BCD methods (e.g. GD or SGD) (Nutini et al., 2015; Zhao et al., 2014). It complicates both optimizer design and convergence analysis. Consequently, combining BCD with Adam is not a straightforward task.
@@ 2.1 Motivation
[…]
In this work, we conduct exploratory experiments to investigate this correlation among minima produced by different LLM fine-tuning methods. We conduct two sets of experiments. First, we fine-tune Pythia-160M on a subset of the FLAN dataset [^3] The subset used is ‘definite_pronoun_resolution_10templates,’ available at https://huggingface.co/datasets/Muennighoff/flan. The learning rate is 2e-5 and the batch size is set as 64. using Adam (Kingma & Ba, 2014) and Lion (Chen et al., 2024). As illustrated in Figure 2(a)(a), Adam and Lion converge to distinct minima. Notably, while both optimizers achieve similar fine-tuning losses, the minimum reached by Adam is much closer to the pre-trained model compared to Lion, being only about 20% of Lion’s distance. (Specifically we measure the model distance by the average $L_{2}$ norm difference over parameter blocks; see Appendix C.4 for the formal definition). As shown in Figure 2(a)(b), Adam exhibits a smaller increase in pre-training loss than Lion. Further, Adam leads to better preservation in general capability than Lion (see Appendix B.2).
Second, we conduct a larger-scale exploratory experiment to test the generality of this observation. We fine-tune a larger model, LLaMA2-7B, on the MetaMathQA dataset (Yu et al., 2024b) using three optimizers: Adam, Lion, and our proposed MoFO (to be introduced in Section 2.2). To achieve varying distances from the pre-trained model, we fine-tune each model for 0.5, 1, 1.5, and 2 epochs (309 steps per epoch). Details are provided in Appendix C.3. Figure 2(b)(a) demonstrates a strong positive correlation between distance and the increase in pre-training loss. Figure 2(b)(b) indicates a negative correlation with MMLU accuracy (Hendrycks et al., 2021), which measures the preservation of factual knowledge, once the model has been trained for at least one epoch. Taken together, these results suggest that, during LLM fine-tuning, remaining closer to the pre-trained state appears to be associated with reduced forgetting. Further discussion of early-training behavior and a comparison of different optimizers are provided in Appendix B.3.
(b) (a) Loss changes on the RedPajama dataset and (b) average accuracy changes on MMLU benchmark (measuring the preservation of factual knowledge) of Llama-2-7B after fine-tuning on MetaMathQA using Adam, Lion, and MoFO for 0.5, 1, 1.5, 2 epochs. We note that RedPajama project was explicitly designed as an open-source reproduction of the LLaMA training dataset (Weber et al., 2024). Thus, it serves as a reasonable proxy for the original LLaMA-2 training dataset since the latter has not been publicly released. See Appendix B.1 for the rationale. The results show a strong positive correlation between the distance from the pre-trained model and the extent of forgetting after one epoch. Further discussion of early-training behavior and a comparison of different optimizers are provided in Appendix B.3.
@@ 2.2 Formulation of MoFO
1: Input: Filtering threshold $\alpha$, number of partitions $B$ with the $k$-th partition of size $d_{k}$, hyperparameters $\beta_{1},\beta_{2}$ of Adam optimizer, learning rate schedule $\{\eta_{t}\}$.
@@
[…]
@@
Figure 2: Illustration of MoFO.
We formally introduce the Momentum-Filtered Optimizer (MoFO) in Algorithm 1. First, all model parameters are partitioned into $B$ blocks. At each iteration, MoFO first computes the gradient and momentum terms for parameters in each block following the standard rule of Adam, as shown in Lines 5-9. Then, MoFO selects and updates the parameter entries with the largest $\alpha$ momentum magnitudes in each parameter block, as shown in Lines 10-13, where the update fraction $\alpha$ is a pre-determined hyperparameter. This momentum filtering mechanism is illustrated in Figure 2.
Mathematically, the filter can be represented as follows. Consider a momentum vector $m=(m^{(1)},\dots,m^{(B)})$, where each $m^{(k)}\in\mathbb{R}^{d_{k}}$ corresponds to the $k$-th block of parameters with dimensionality $d_{k}$. The top-$\alpha$ filter, denoted as $\texttt{FLT}_{\alpha}(m)$, is defined as $\texttt{FLT}_{\alpha}(m)=(\texttt{FLT}_{\alpha}^{(1)}(m),\dots,\texttt{FLT}_{\alpha}^{(B)}(m))$, where the $i$-th entry of $\texttt{FLT}_{\alpha}^{(k)}(m)$ is given by
[…]
For the parameter partitioning, we note that the network architecture is naturally composed of different modules (e.g., weight matrices, and bias terms). In the PyTorch implementation, the parameters of different modules (along with their gradients and momenta) are naturally stored in separate data tensors. Therefore, we adopt the default partitioning of model parameters as implemented in PyTorch. For Transformers, this means that parameters such as query (Q), key (K), value (V) weights in the attention layers, as well as feed-forward network (FFN) weights, are grouped into distinct partitions following PyTorch’s default scheme. This allows us to select and update the top-$\alpha$ parameters in each block without introducing much implementation overhead. See Appendix C.4 for further explanation of the partitioning.
@@ 3.1 Convergence Result
Suppose that the first- and second-order momentum hyperparameters $\beta_{1}$ and $\beta_{2}$ satisfy $0<\beta_{1}<\sqrt{\beta_{2}}<1$. The learning rate schedule at step $t$ is $\eta_{t}=\eta/\sqrt{t}$ for some $\eta>0$. Then, under Assumption 1, MoFO satisfies
@@
$\min_{0\leq t\leq T-1}\|\nabla\mathcal{L}(\theta_{t})\odot\texttt{FLT}_{\alpha}\big(\nabla\mathcal{L}(\theta_{t})\big)\|_{1}=\mathcal{O}\left(\frac{\log T}{\sqrt{T}}\right)\quad\text{as }T\to\infty.$
Moreover, this bound directly implies
@@
$\min_{0\leq t\leq T-1}\|\nabla\mathcal{L}(\theta_{t})\|_{p}=\mathcal{O}\left(\frac{\log T}{\sqrt{T}}\right)\quad\text{as }T\to\infty,$
for any $p\in[1,\infty]$.
Although MoFO is designed to mitigate forgetting by updating only a small subset of parameters at each step, it is guaranteed to converge to a critical point of the fine-tuning loss function under the Lipschitz smoothness assumption. This result provides theoretical evidence that MoFO can achieve competitive performance in fine-tuning tasks.
Proof Sketch of Theorem 1:.
Our proof is inspired by the convergence analysis for full-batch Adam in Shi et al. (2021), but we introduce additional techniques tailored to MoFO’s filtering mechanism. We will highlight these additions precisely at the points where they arise below.
Let $g_{t}=\nabla\mathcal{L}(\theta_{t-1})$. A central step is to establish, for suitable constants $C_{1},C_{2}>0$,
$\frac{C_{1}}{\sqrt{t}}\|g_{t}\|\leq\mathcal{L}(\theta_{t-1})-\mathcal{L}(\theta_{t})+\frac{C_{2}}{t},$ (1)
Summing this inequality from $t=1$ to $T$ and using $\sum_{t=1}^{T}t^{-1}=\log T+\mathcal{O}(1)$ yields the convergence result for Adam in terms of a diminishing norm of gradient, given by
$\min_{1\leq t\leq T}\|g_{t}\|=\mathcal{O}\left(\frac{\log T}{\sqrt{T}}\right).$ (2)
Choice of norm and two subgoals.
In finite-dimensional spaces, all norms are equivalent, so convergence statements like (1)–(2) can be expressed in any fixed norm up to norm-equivalence constants. In practice, however, specific analyses instantiate (1) with a particular norm: Shi et al. (2021) work with the $L_{1}$ norm $\|g_{t}\|_{1}$ for full-batch Adam, whereas our full-batch MoFO analysis will use the $L_{1,\text{top-}\alpha}$ norm $\|g_{t}\odot\texttt{FLT}_{\alpha}(g_{t})\|_{1}$, which will be defined in Appendix A.1.
To keep the logic precise, we separate our argument into two subgoals:
Step I (Key inequality for MoFO). Show that there exist constants $C_{1},C_{2}>0$ (independent of $t$) such that
$\frac{C_{1}}{\sqrt{t}}\,\big\|g_{t}\odot\texttt{FLT}_{\alpha}(g_{t})\big\|_{1}\;\leq\;\mathcal{L}(\theta_{t-1})-\mathcal{L}(\theta_{t})+\frac{C_{2}}{t}.$ (3)
Here, we recall that $\texttt{FLT}_{\alpha}(\cdot)$ preserves the $\alpha$-fraction of largest-magnitude coordinates in each partition and zeros out the rest.
Step II (Norm property of the left-hand side of (3)). This component is one of our new technical ingredients (absent from Shi et al. (2021)). We first verify that the mapping $x\mapsto\|x\odot\texttt{FLT}_{\alpha}(x)\|_{1}$ defines a norm on $\mathbb{R}^{d}$, which is referred to as $L_{1,\text{top-}\alpha}$ norm. This is proved in Proposition 1 (Appendix A.1) by checking nonnegativity and definiteness, positive homogeneity, and the triangle inequality. Moreover, for the $L_{p}$ upper bound in Theorem 1, we use the norm equivalence between the $L_{1,\text{top-}\alpha}$ norm and the $L_{p}$ norm, as shown in Lemma 2.
With Step II established, it remains to prove the key inequality in Step I. A direct adaptation of Shi et al. (2021) to MoFO is infeasible due to structural differences. We proceed as follows:
  1. (i)
    Recap key elements from Shi et al. (2021);
  2. (ii)
    Identify challenges in extending them to MoFO;
  3. (iii)
    Resolve these challenges by carefully handling the momentum filter.
Part (i): Key elements of Shi et al. (2021). For full-batch Adam with bias-corrected moments $\hat{m}_{t},\hat{v}_{t}$ and the learning rate schedule $\eta_{t}=\eta/\sqrt{t}$, the parameter update is
$\theta_{t}-\theta_{t-1}=-\frac{\eta}{\sqrt{t}}\cdot\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}}.$
By the $L$-smoothness of the loss and the descent lemma,
$\frac{\eta}{\sqrt{t}}\sum_{i=1}^{d}g_{i,t}\frac{\hat{m}_{i,t}}{\sqrt{\hat{v}_{i,t}}}\leq\mathcal{L}(\theta_{t-1})-\mathcal{L}(\theta_{t})+\frac{L}{2}\|\theta_{t}-\theta_{t-1}\|_{2}^{2}.$ (4)
Lemma 5 in Appendix A.2 lower-bounds the per-coordinate contribution as
$g_{i,t}\frac{\hat{m}_{i,t}}{\sqrt{\hat{v}_{i,t}}}\geq A|g_{i,t}|-\frac{B}{\sqrt{t}},$
for some constants $A,B>0$. Substituting this into (4), summing over coordinates, and controlling the quadratic term in (4) yields the fundamental inequality (1) with the $L_{1}$-norm:
$\frac{C_{1}}{\sqrt{t}}\|g_{t}\|_{1}\leq\mathcal{L}(\theta_{t-1})-\mathcal{L}(\theta_{t})+\frac{C_{2}}{t},$
Part (ii): Challenges in extending to MoFO. For the full-batch version of MoFO, the parameter update becomes
$\theta_{t}-\theta_{t-1}=-\frac{\eta}{\sqrt{t}}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}}\odot\texttt{FLT}_{\alpha}(m_{t}),$
i.e., the step is filtered by momentum magnitudes (NOT gradient magnitudes). Building upon the convergence analysis of Shi et al. (2021) in Part (i) leads to
$\frac{C_{1}}{\sqrt{t}}\|g_{t}\odot\texttt{FLT}_{\alpha}(m_{t})\|_{1}\leq\mathcal{L}(\theta_{t-1})-\mathcal{L}(\theta_{t})+\frac{C_{2}}{t},$ (5)
which brings a notable difference to the target key inequality (3): the inequality (5) applies the momentum filter $\texttt{FLT}_{\alpha}(m_{t})$, whereas the desired bound applies the gradient filter $\texttt{FLT}_{\alpha}(g_{t})$. This introduces a non-trivial challenge because the $\|g_{t}\odot\texttt{FLT}_{\alpha}(m_{t})\|_{1}$ being small does not naturally imply that $\|g_{t}\odot\texttt{FLT}_{\alpha}(g_{t})\|_{1}$ is small: large entries of $g_{t}$ might be excluded if their momenta lags. Consequently, additional analysis is required to bound the discrepancy between them.
Part (iii): How we overcome the challenge. Here we introduce another new technique to address the discrepancy between $\|g_{t}\odot\texttt{FLT}_{\alpha}(m_{t})\|_{1}$ and $\|g_{t}\odot\texttt{FLT}_{\alpha}(g_{t})\|_{1}$. We deal with the challenge via Lemma 1 in Appendix A.1, which establishes that the $L_{1}$-deviation in filtered outputs is Lipschitz-stable under input perturbations. Specifically, for any $x,y\in\mathbb{R}^{d}$:
$\underbrace{\left\|x\odot\texttt{FLT}_{\alpha}(x)\right\|_{1}-\left\|x\odot\texttt{FLT}_{\alpha}(y)\right\|_{1}}_{\text{top-}\alpha\text{ filtered error}}\leq 2\underbrace{\|x-y\|_{1}}_{\text{input error}}.$
This result points out that while the filter $\texttt{FLT}_{\alpha}(\cdot)$ itself is unstable under perturbation, the filtered output remains controllable through. It effectively "smooths" the discontinuity that would plausibly prevent convergence analysis.
With the Lipschitz stability established, we now deal with the challenge. We first control the $L_{1}$ distance between the bias-corrected momentum $\hat{m}_{t}=\frac{m_{t}}{1-\beta_{1}^{t}}$ and the gradient $g_{t}$ by showing that $\|\hat{m}_{t}-g_{t}\|_{1}=\mathcal{O}(1/\sqrt{t})$. Second, we apply Lemma 1 with $x=g_{t}$ and $y=\hat{m}_{t}$ and yield
$\left\|g_{t}\odot\texttt{FLT}_{\alpha}(g_{t})\right\|_{1}-\left\|g_{t}\odot\texttt{FLT}_{\alpha}(\hat{m}_{t})\right\|_{1}\leq 2\|\hat{m}_{t}-g_{t}\|_{1}=\mathcal{O}(1/\sqrt{t}).$
Since the filtering function $\texttt{FLT}_{\alpha}(\cdot)$ is invariant under positive scaling, we can definitely replace $\hat{m}_{t}$ with the original momentum $m_{t}$:
$\left\|g_{t}\odot\texttt{FLT}_{\alpha}(g_{t})\right\|_{1}-\left\|g_{t}\odot\texttt{FLT}_{\alpha}(m_{t})\right\|_{1}\leq 2\|\hat{m}_{t}-g_{t}\|_{1}=\mathcal{O}(1/\sqrt{t}).$
Combining it with (5) and subsuming residual $\mathcal{O}(1/t)$ terms, we obtain our target key inequality (3):
@@
$\frac{C_{1}}{\sqrt{t}}\left\|g_{t}\odot\texttt{FLT}_{\alpha}(g_{t})\right\|_{1}\leq\mathcal{L}(\theta_{t-1})-\mathcal{L}(\theta_{t})+\frac{C_{2}}{t}.$
In conclusion, our statement of Step I primarily comprises:
  • Step I.1 (Part (i) and (ii)). Extending the Adam convergence framework (Shi et al., 2021) to incorporate MoFO’s momentum filtering mechanism, yielding (5);
  • Step I.2 (Part (iii)). Resolving the momentum-gradient filter discrepancy via our Lipschitz stability analysis (Lemma 1), which is a new technique introduced in this work.
Combined with the norm property in Step II—another new ingredient not present in Shi et al. (2021)—, these results enable us to establish Theorem 1.